### v3

# Analyzing User Characteristics of Hate Speech Spreaders on Social Media

Abstract.
Hate speech on social media threatens the mental and physical well-being of individuals and contributes to real-world violence. Resharing is an important driver behind the spread of hate speech on social media. Yet, little is known about who reshares hate speech and what their characteristics are. In this paper, we analyze the role of user characteristics in hate speech resharing across different types of hate speech (e.g., political hate). For this, we first cluster hate speech posts using large language models into different types of hate speech. Then we model the effects of user attributes on users’ probability to reshare hate speech using an explainable machine learning model. To do so, we apply debiasing to control for selection bias in our observational social media data and further control for the latent vulnerability of users to hate speech. We find that, all else equal, users with fewer followers, fewer friends, fewer posts, and older accounts share more hate speech. This shows that users with little social influence tend to share more hate speech. Further, we find substantial heterogeneity across different types of hate speech. For example, racist and misogynistic hate is spread mostly by users with little social influence. In contrast, political anti-Trump and anti-right-wing hate is reshared by users with larger social influence. Overall, understanding the factors that drive users to share hate speech is crucial for detecting individuals at risk of engaging in harmful behavior and for designing effective mitigation strategies. Disclaimer: This work contains terms that are offensive and hateful. to the nature of the work.
hate speech, social media, propensity score, recommender systems, user characteristics, resharing
copyright: noneccs: Social and professional topics User characteristicsccs: Human-centered computing Social content sharingccs: Human-centered computing Social mediaccs: Applied computing Sociology
1. Introduction
Online hate speech poses a significant problem to society. Around one in four U.S. adults has experienced online hate (Vogels, 2021), and, in particular, young adults suffer from the psychological consequences (Saha et al., 2019). Hate speech not only threatens users’ mental and physical well-being but also has the potential to fuel radicalization and incite real-world violence (Müller and Schwarz, 2021), as was the case for the suspect of the 2018 Pittsburgh synagogue shooting (Roose, 2018).
On social media, reshares are vital in disseminating content beyond immediate circles (Subbian et al., 2017). One risk of reshares is that harmful content can go viral, such as propaganda (Geissler et al., 2023), fake news (Cheng et al., 2021; Shao et al., 2018; Pröllochs et al., 2021a, b), and content from untrustworthy sources (Guess et al., 2023). Prior research found that reshares also make hate speech go disproportionally more viral than normal content (Maarouf et al., 2024; Mathew et al., 2019), which is exploited by hateful users to reach more people (Goel et al., 2023).
However, prior work has mostly looked into who posts hate speech (Hua et al., 2020; Ribeiro et al., 2018) and who is targeted by it (ElSherief et al., 2018). Hateful users are characterized by high activity levels, complex word usage, and high connectivity among each other (Hua et al., 2020; Ribeiro et al., 2018). Similarly, targets of hate speech show high levels of activity, but they also tend to be older accounts and have more followers (ElSherief et al., 2018). However, no work has been conducted to profile the users who reshare hate speech, although reshares help hate speech reach larger audiences and are an important driver to the virality of hate speech (Maarouf et al., 2024). Hence, in this work, we analyze the effects of attributes (such as, e.g., verified status, account age, or number of posts, followers, and friends) of users who reshare hate speech while taking into account the different types of hate speech.
Estimating the effects of user attributes on users’ probability to reshare hate speech is difficult. One challenge is the inherent selection bias in observational social media click data, which comes from the users’ self-selection of which content to interact with as well as the recommendations from online platforms (Schnabel et al., 2016). For example, users only interact with content that they choose and online platforms only recommend content that they assume a user is likely to engage with. We later employ inverse propensity scoring (IPS) (Rosenbaum and Rubin, 1983) to reweight the observed interactions between users and hate speech posts for debiasing. Another reason is that we have to control for the varying vulnerability of users to hate speech as that can cause bias in the estimates of the effects being studied. Previous research argued that past latent vulnerability (PLV) of a user to specific content influences both the user attributes and the probability to reshare (Benevenuto et al., 2009; Cheng et al., 2021). Users may have varying PLV as some users simply share more hate speech than others. Also, the risk of exposure to hate speech is different for each user, e.g. how many times a user has seen hate speech in their feed. In our setting, we control for PLV to hate speech.
In this work, we analyze the role of user characteristics in hate speech resharing across different types of hate speech. Our framework follows three steps: 1 We first cluster hate speech posts into their type using BERTopic (Grootendorst, 2022) and employ automatic cluster labeling with LLAMA-3 (Meta, 2024). This way, we obtain a topic label for each hate speech post. 2 We then model the past latent vulnerability of each user to hate speech. For this, we debias the observational data by adapting inverse propensity scoring (IPS) from causal inference (Rosenbaum and Rubin, 1983). This reweights the likelihood of users to be exposed to hate speech (Liang et al., 2016; Schnabel et al., 2016) and we obtain debiased propensity scores. We employ the debiased propensity scores to model the debiased PLV of users to hate speech by learning latent representations in the form of debiased embeddings. 3 We then model the effects of user attributes on the probability to reshare hate speech while controlling for the PLV. Here, we compute the likelihood to reshare each type of hate speech per user by looking at past reshares. Then, we model the (non-linear) effects of the user attributes (verified status, account age, number of posts, followers, and friends) on the user’s probability to reshare hate speech in the future using an explainable boosting machine (EBM) (Nori et al., 2019). [^1] Codes and data for reproducibility are available via an anonymous repository for review and will be available in a public repository upon publication: https://anonymous.4open.science/r/analyzing_user_characteristics_hatespeech_spreaders-2042/README.md
2. Related Work
2.1. Hate speech on social media
[…]
One stream of literature conducted content-level analyses. Here a frequent theme is to develop machine learning models that can detect and subsequentially remove hate speech. Such systems aim to detect hate speech, for example by using large language models (Barbieri et al., 2020; Djuric et al., 2015; Mozafari et al., 2019) or other natural language processing methods (Badjatiya et al., 2017; Davidson et al., 2017).
Another stream of literature conducted user-level analyses. Targets of hate speech tend to be more popular or high-profile users (ElSherief et al., 2018), while users that post hate speech are characterized by higher social media activity and tend to act in more connected networks (Hua et al., 2020; Ribeiro et al., 2018). When looking at hate speech in online political discourse, people of color of the Democratic Party, white Republicans, and women in general tend to be the main targets (Solovev and Pröllochs, 2022).
However, none of the above works analyze resharing of hate speech. Yet, reshares are what makes content go viral and play a vital role in the dissemination of hate speech online. Moreover, no work has been conducted that analyzes who reshares hate speech and specifically which type of hate speech. This is in contrast to our work, which aims to answer which user attributes make users reshare hate speech and specifically the different types of hate speech.
2.2. Inverse propensity scoring
Click data from observational social media data are subject to selection bias. Specifically, previous research identifies two sources of selection bias: self-selection by the users and recommendations from online platforms (Schnabel et al., 2016). On the one hand, users self-select the content they interact with on social media. On the other hand, online platforms tend to recommend only content to users, that a user is likely interested in. This means that click data are biased because the observed interactions were influenced by self-selection and recommendation bias. Hence, we must carefully account for selection bias in our observational click data.
To alleviate selection bias in observational click data, previous works draw upon a popular technique from causal inference, specifically inverse propensity scoring (IPS) (Rosenbaum and Rubin, 1983; Imbens and Rubin, 2015; Rubin, 2001). Each hate speech post is reweighted with its inverse propensity score, i.e. the probability that a user is exposed to the post. This way, we transform the observational click data as though it came from a randomized trial where users are randomly shown items (Liang et al., 2016; Schnabel et al., 2016). By doing so, the selection bias is alleviated and debiased propensity scores are obtained. Some works have applied IPS to learn and evaluate debiased recommendation models (Liang et al., 2016; Schnabel et al., 2016). Others have applied IPS to debias observational data used to learn users’ fake news resharing behavior on social media (Cheng et al., 2021). In particular, the estimated propensity scores of fake news were debiased in three ways: (a) by using the popularity of shared fake news items; (b) by using the popularity of resharing users; or (c) by employing a neural network (Cheng et al., 2021).
In the context of our work, the propensity scores describe the probability that a user is exposed to hate speech. Here, we also deal with selection bias of our observational click data. Hence, we later also adopt the debiased propensity score estimations and integrate them into our framework for modeling hate speech resharing. This way, we obtain debiased propensity scores for the exposure of users to hate speech, which we can use to model debiased, latent representations of the vulnerability to hate speech.
2.3. Modeling past latent vulnerability
In general, the vulnerability of users to specific online content cannot be directly observed from data but instead can be represented by latent embeddings. Previous work modeled PLV through recommendation algorithms (Bonner and Vasile, 2018). Recommendation algorithms take into account previous interactions of users with online content in order to suggest new content the user is likely to interact with. The recommendation algorithms are then evaluated by the coherence of recommendation candidates with past user behavior. This way, they inherently mimic user behavior (Bonner and Vasile, 2018). For example, prior work leveraged recommendation algorithms to model the PLV of users to fake news, which was used as a confounder to find the attributes that cause users to reshare fake news online (Cheng et al., 2021). However, to the best of our knowledge, no work modeled the PLV of users to hate speech on social media.
We later transfer the idea of recommendation algorithms to model the PLV of users to hate speech. In particular, we employ recommendation algorithms to learn the debiased PLV of users to hate speech from IPS-reweighted data. By doing so, we represent the PLV as a debiased embedding.
3. Methodology
3.1. Overview
Our objective is to model the effects of user attributes $A$ on the users’ probability to reshare different types of hate speech $Y$ while accounting for the user’s (latent) vulnerability to hate speech $X$. For this, we use BERTopic to cluster posts into different types of hate speech and employ a state-of-the-art, explainable debiasing framework to model the effects.
Input: We input the interactions of users with hate speech posts as a bipartite graph into our framework. Let $\mathcal{U}=\{u_{1},\ldots,u_{i}\}$ denote users and =$\mathcal{H}=\{h_{1},\ldots,h_{j}\}$ denote the hate speech posts. The interaction between user $u\in\mathcal{U}$ and hate speech post $h\in\mathcal{H}$ is then a binary variable denoted as $S_{uh}=\{0,1\}$, where $S_{uh}=1$ if user $u$ shared hate speech post $h$, otherwise $S_{uh}=0$.
1
Clustering hate speech: In the first step, we employ BERTopic (Grootendorst, 2022) combined with HDBSCAN to cluster hate speech posts $\mathcal{H}$. Subsequently, we label the clusters using automatic labeling with LLAMA-3 (Meta, 2024). Hence, we generate clusters of types of hate speech $\mathcal{C}=\{c_{1},\ldots,c_{n}\}$.
2
Modeling past latent vulnerability: In the next step, we model the past latent vulnerability $X$ of users to hate speech. For this, we first apply IPS reweighting to debias our observational data, which reweights the propensity $\pi$ of hate speech posts to mimic a randomized trial, and, thereby, we address the issue of selection bias from above. Then, we model the past latent vulnerability to hate speech $X_{u}$ for each user $u$.
3
Modeling hate speech resharing: In the final step, we model the effects of user attributes $A$ on users’ probability to reshare hate speech $Y$ while controlling for PLV ${X}$. We write the individual user’s probability to reshare hate speech in general as outcome variable $Y_{u}\in[0,1]$ and the probability to reshare hate speech type $c$ as $Y_{uc}\in[0,1]$. Then, we use an explainable boosting machine (EBM) (Nori et al., 2019) to predict the probability to reshare hate speech. For each user $u$ with $m$ user attributes denoted as vector ${a}_{u}=(a_{1},\ldots,a_{m})$, ${a}_{u}\in A$ and PLV embedding ${X}_{u}$ we model the probability to reshare hate speech $Y_{u}$ ($Y_{uc}$ for different types of hate speech $c$).
3.2. Clustering hate speech (Step 1)
Given hate speech posts $\mathcal{H}$, we employ BERTopic (Grootendorst, 2022), which leverages BERT embeddings and topic modeling, combined with HDBSCAN for clustering. First, we remove stop words, URLs, and emojis. Then, we generate embeddings for each hate speech post using the pre-trained BERT model all-MiniLM-L6-v2 to capture the semantic content. Next, we reduce the embedding dimensionality using UMAP to enhance performance and interpretability. Lastly, we apply HDBSCAN to the reduced embeddings to identify clusters $\mathcal{C}$. HDBSCAN is highly flexible as it does not require pre-specified numbers of clusters and it is suitable for noisy data, which makes it an ideal choice for social media data. In addition, HDBSCAN can deal well with clusters that vary in shape and size (Campello, Ricardo J. G. B. et al., 2013).
Subsequently, we label the clusters using automatic labeling using LLAMA-3 model Llama-3-70b-chat-hf (Meta, 2024). We generate five names per cluster and manually pick the most fitting one by analyzing representative posts. We hereby obtain concise summaries of the primary themes within each cluster, thus providing clear and contextually relevant labels.
3.3. Modeling past latent vulnerability (Step 2)
3.3.1. IPS reweighting
Given the users $\mathcal{U}$, the hate speech posts $\mathcal{H}$, and the user-hate speech interactions $S_{uh}$, we apply IPS reweighting to alleviate selection bias and obtain debiased propensity scores $\pi$. The propensity score describes the probability that a user is exposed to hate speech.
Assumption: As in (Joachims et al., 2017), we assume that users share posts when (a) they have been exposed to the post, denoted as binary variable $E_{uh}\in\{0,1\}$ ($E_{uh}=1$ if user $u$ was exposed to hate speech post $h$, $E_{uh}=0$, otherwise), and (b) when they are interested in the post, denoted as binary variable $I_{uh}\in\{0,1\}$ ($I_{uh}=1$ if the user is interested in the post, $I_{uh}=0$, otherwise). We further assume that resharing of hate speech is missing-not-at-random (MNAR) (Little and Rubin, 2020). Hence, the probability of user $u\in\mathcal{U}$ spreading a hate speech post $h\in\mathcal{H}$ is a product of the probability of exposure and the probability of interestingness of a post (Saito, 2020). Formally, we have
(1) $\begin{split}P(S_{uh}=1)&=\underbrace{P(E_{uh}=1)}_{\text{exposure prob.}}\quad\cdot\underbrace{P(I_{uh}=1)}_{\text{interestingness prob.}}\\ &=\theta_{uh}\cdot\iota_{uh}\end{split}$
for all $S_{uh}\in S$ and where $\theta_{uh}>0$ is the probability that user $u$ is exposed to hate speech post $h$ and $\iota_{uh}>0$ is the probability that user $u$ is interested in hate speech post $h$. Hence, the question arises how to compute $\theta_{uh}$. Our idea in the following is to leverage the propensity score $\pi$ for that, where the propensity score is defined as $\pi=P(E_{uh}=1)$. However, the propensity score gives a biased estimate of $\theta_{uh}$ due to the observational nature of social media data, and, instead, we then use a debiased propensity score.
Why the propensity score is a biased estimate of $\theta_{uh}$: Simply computing $\pi$ from our data gives a biased estimate for $\theta_{uh}$. The reason is that, when working with observational click data, we face two major challenges: First, we only record positive interactions between users and hate speech posts (i.e., hate speech posts that are reshared by the user). Negative interactions (i.e., hate speech posts that were not reshared by the users) are not recorded. Therefore, we do not know whether (a) the user was not interested in a given post or whether (b) the user was not exposed to the post. Second, the observed interactions are likely subject to selection bias. Users are more likely to reshare hate speech posts that are more prevalent, and online platforms are more likely to recommend popular posts to users. Because of the above, the propensity scores $\pi$ is biased when computed from observational social media data, i.e.,
(2) $\pi_{\mathrm{biased}}=P(S_{uh}=1\,\mid\,I_{uh}=1)\neq\theta_{uh}$
Debiasing the propensity score: Here, we propose to leverage IPS to learn debiased propensity scores based on the observed social media data. IPS acts as a reweighting mechanism that assigns larger weights to hate speech posts that are less likely to be observed, thereby creating a pseudo-randomized trial where users are shown posts at random. By doing so, we will later be able to learn debiased embeddings of the past latent vulnerability to hate speech. Following earlier research (Cheng et al., 2021), we use three different estimation procedures to obtain unbiased propensity scores:
  1. (1)
    Virality-based propensity score: This propensity score takes into account the virality of a hate speech post $h$, i.e.,
    (3) $\pi_{h}^{\mathrm{virality}}=\hat{\theta}^{\mathrm{virality}}_{h}=\left(\frac{\sum_{u\in\mathcal{U}}S_{uh}}{\max_{h\in\mathcal{H}}\sum_{u\in\mathcal{U}}S_{uh}}\right)^{\mu},$
    where we assume that the probability that a user $u$ is exposed to a hate speech post $h$ is higher if the post has gone more viral (as measured by the number of reshares in the bipartite graph). The IPS for rare events could be very high. Hence, we apply a smoothing operator $\mu=0.5$ for more stable propensity scores (see Appendix D for a robustness check with varying $\mu$).
  2. (2)
    Follower-based propensity score: This propensity score takes into account the virality of the hate speech post $h$ as well as the popularity of the resharing user $u$ (as measured as the number of followers $F_{u}$). Formally, it is given by
    (4) $\pi_{h}^{\mathrm{follower}}=\hat{\theta}^{\mathrm{follower}}_{uh}=\left(\frac{\sum_{u\in\mathcal{U}}S_{uh}\cdot F_{u}}{max_{h\in\mathcal{H}}\sum_{u\in\mathcal{U}}S_{uh}\cdot F_{u}}\right)^{\mu},$
    with $\mu=0.5$. Here, we also account for the bias induced by the popularity of the user, since popular users are more likely to be exposed to hate speech (ElSherief et al., 2018).
  3. (3)
    Neural network-based propensity score: This propensity score is estimated by a neural network via
    (5) $\pi_{h}^{\mathrm{neural}}=\hat{\theta}^{\mathrm{neural}}_{h}=\sigma({e_{h}}),$
    where ${e_{h}}$ is the content of a given hate speech post as learned by LDA topic modeling and $\sigma()$ is the sigmoid function. This way, we implicitly encode the hate speech post’s content in the latent representation.
As a result, we obtain debiased propensity scores. Later, we perform an ablation study where we compare the three variants (1)–(3) of our recommendation algorithm to the biased algorithm and find that the debiased recommendation algorithms perform better.
3.3.2. Modeling past latent vulnerability
To model the past latent vulnerability (PLV) of users to hate speech (which gives our control variable ${X}$), we build upon concepts from recommendation systems (Bonner and Vasile, 2018). Thereby, we address one particular difficulty in that both the interestingness $I_{uh}$ and the exposure $E_{uh}$ are not observable. Rather, only the interactions $S_{uh}$ between users and hate speech posts are observable on social media platforms. Hence, we learn a recommender algorithm that ranks content according to the interestingness of users simply using the observed interactions.
Recommender algorithm: In our framework, we employ a Bayesian personalized ranking matrix factorization (BPRMF) (Rendle et al., 2009). We optimize the parameters in the BPRMF using a pairwise Bayesian personalized ranking (BPR) loss (Rendle et al., 2009). By using BPRMF, we address the challenges of missing negative interactions, since it assumes that the observed interactions better explain users’ preferences than unobserved ones and should hence be assigned a higher predictive score. Since we employ three different ways to debias propensity scores, we also obtain three different variants of our debiased recommendation model: BPRMF-V (using $\pi_{\mathrm{virality}}$); BPRMF-F (using $\pi_{\mathrm{follower}}$); and BPRMF-NN (using $\pi_{\mathrm{neural}}$). In Appendix D we perform a robustness check where we compare the performance of our unbiased recommendation models and the biased recommender. We find that the recommendation algorithms with debiased propensity scores are better at predicting the users vulnerability to hate speech. For each user $u$, we model PLV to hate speech as a 64-dimensional embedding ${X}_{u}$. For comparison, the base model BPRMF is biased since it uses $\pi_{\mathrm{biased}}$ as the propensity score.
[…]
(6) $L_{\mathrm{ideal}}(\hat{R})=\frac{1}{|\mathcal{D}|}\sum_{(u,h,g)\in\mathcal{D}}\iota_{uh}\large(1-\iota_{ug}\large)\,\ell(\hat{R}_{uhg}),$
where $\ell=-\mathrm{ln}\sigma(\cdot)$ is the local loss for the triplet $(u,h,g)$, $\iota_{uh}$ is the interestingness of the post $h$ for user $u$, and $\hat{R}_{uhg}$ is the difference between the predicted scores for hate speech posts $h$ and $g$. However, the interestingness is unobservable and only the interactions $S_{uh}$ are available. BPR then assumes that interacted messages should be ranked higher than all the other non-interacted messages and optimizes the following loss function to obtain latent factors:
(7) $L_{\mathrm{BPR}}(\hat{R})=\frac{1}{|\mathcal{D}|}\sum_{(u,h,g)\in\mathcal{D}}S_{uh}\large(1-S_{ug}\large)\,\ell(\hat{R}_{uhg}).$
However, the loss function of the BPR model is biased towards the ideal pairwise loss, because BPR treats all non-interacted messages as if the user was not interested in them and does not deal with messages the user was not exposed to (Saito, 2020). To deal with the missing data in implicit feedback, we follow previous research (Cheng et al., 2021) and employ propensity scores as $\theta_{uh}$ for the unbiased estimator of the ideal pairwise loss as follows
(8) $L_{\mathrm{unbiased}}(\hat{R})=\frac{1}{|\mathcal{D}|}\sum_{(u,h,g)\in\mathcal{D}}\frac{S_{uh}}{\theta_{uh}}\large(1-\frac{S_{ug}}{\theta_{ug}}\large)\,\ell(\hat{R}_{uhg}).$
Of note, propensity scores can often become extremely small, e.g., hate speech posts with low virality also have low exposure probability. This leads to large variances in IPS-based approaches. We reduce the variance by employing a non-negative loss (Saito, 2020) given by
(9) $\hat{L}_{\textrm{non-neg}}(\hat{R})=\frac{1}{|\mathcal{D}|}\sum_{(u,h,g)\in\mathcal{D}}\max\{\ell(\hat{R}_{uhg}),0\}.$
In addition, we add a $l_{2}$-regularization for the user embeddings and post embeddings, ${U}$ and ${H}$, respectively. The final loss function is
(10) $\operatorname*{arg\,min}_{{U},{H}}\hat{L}_{\textrm{non-neg}}(\hat{R})+\lambda\large(||{U}||_{2}^{2}+||{H}||_{2}^{2}\large),$
where $\lambda$ is the weight of the $l_{2}$-regularization for the latent representations.
3.4. Modeling hate speech resharing (Step 3)
3.4.1. Likelihood to reshare hate speech
For each user $u$, we now compute the probability to reshare hate speech in general $Y_{u}$ as well as the probability to reshare different types of hate speech $Y_{uc}$. Formally, we define the user’s probability to reshare hate speech $Y_{u}\in[0,1]$ as
(11) $Y_{u}=\frac{n^{u}_{\mathrm{hate}}}{n^{u}_{\mathrm{hate}}+n^{u}_{\mathrm{normal}}},$
where $n^{u}_{\mathrm{hate}}$ and $n^{u}_{\mathrm{normal}}$ are the number of hate speech posts and normal posts that user $u$ shared, respectively.
Similarly, the probability to reshare a specific type of hate speech $Y_{uc}\in[0,1]$ is defined as
(12) $Y_{uc}=\frac{n^{u}_{c}}{n^{u}_{\mathrm{hate}}+n^{u}_{\mathrm{normal}}},$
where $n^{u}_{c}$ is the number of hate speech posts that user $u$ has shared from cluster $c$.
3.4.2. Effect estimation
We now model the relationships between the user attributes $A$ and the users’ probability to reshare hate speech $Y$ while controlling for PLV ${X}$ for each user $u$. We model the effects using an explainable boosting machine (EBM) (Nori et al., 2019). EBM is a tree-based, cyclic gradient boosting GAM (generalized additive model) with automatic interaction detection. It provides state-of-the-art performance while being fully explainable.
In our work, we are interested in understanding the effects of the following user attributes $a\in A$: verified ($=1$ if the user is verified, 00 otherwise), account age (in days), #posts (number of posts the user did), #followers, and #friends. Formally, we model
(13) $g(E[Y_{u}])=\beta_{0}+\sum f_{m}(a_{m})+\sum f({X}_{u}),$
where $g$ is the link function that adapts the GAM to the regression task. For simplicity, we use $Y_{u}$ in above equation, which can be replaced by $Y_{uc}$ for each hate speech cluster $c\in C$. By using EBM instead of traditional, linear regression (Cheng et al., 2021), we are able to model the non-linear effects of each user attribute on the users’ probability to reshare hate speech. Still, our results are explainable in that we can interpret the estimated effects.
4. Experimental Setup
4.1. Data
To understand who shares hate speech and which type, we apply our framework to a comprehensive, state-of-the-art dataset (Founta et al., 2018). The dataset consists of normal and hate speech posts from the social media platform X (formerly Twitter) that were human-labeled following the definitions and procedures described in (Founta et al., 2018). For all posts in the original dataset, we then additionally collected the corresponding reshares. As a result, our augmented dataset consists of 25,196 root posts (96.05% normal and 3.95% hateful posts) from 31,905 users. The root posts received $N=$ 65,920 reshares by 61,961 users, out of which 43,929 (66.64%) were reshares of normal root posts and 21,991 (33.36%) were reshares of hateful root posts. As such, hateful root posts were reshared disproportionally often.
Model variables: We analyze the effects of five user attributes on the probability to reshare: $\boldsymbol{a}_{1}$: verified ($=1$ if the user is verified, 0 otherwise); $\boldsymbol{a}_{2}$: account age (in days); $\boldsymbol{a}_{3}$: #post (number of posts by the user); $\boldsymbol{a}_{4}$: #follower (number of followers of the user); and $\boldsymbol{a}_{5}$: #friend (number of friends of the user). The distributions of #post, #follower, and #friend are highly right-skewed, because of which we log these user attributes. Appendix Table 3 shows summary statistics for the different user attributes.
Ethical statement: Data collection was conducted by following standards for ethical research (Rivers and Lewis, 2014). Our analyses are based on publicly available data, and we only report aggregated results. We respect users’ privacy by not publishing usernames and do not attempt to track or de-identify users who decided to remain anonymous. We have received ethical approval by our ethics committee (=IRB) at the University of (anonymized). The approval number is (anonymized).
4.2. Baselines
In this work, we leverage a debiasing framework to model the effects of user attributes on users’ probability to reshare hate speech using an EBM. We refer to our framework as DF-EBM. Note that, for benchmarking, our setting requires baselines that provide explainable results since we aim to provide an understanding of the user attributes that make users reshare hate speech. To the best of our knowledge, we are aware of only one benchmark where the effects of user attributes on the probability to reshare content were modeled: (1) DF-linear (Cheng et al., 2021) performs debiasing but, in contrast to our framework, is fully linear. Formally, instead of allowing for non-linear relationships in step  3, DF-linear models linear effects of user attributes on the probability to reshare content. It is thus explainable as it outputs coefficients for each user attribute, but it cannot capture non-linearities and is thus not flexible since the modeled effects are assumed to be linear.
In addition, we design the following variants of our debiasing framework that we also use as baselines: (2) DF-ExNN where we replace the EBM with an explainable neural network (Yang et al., 2020). The explainable neural network enhances the explainability of a vanilla neural network by adding architecture constraints, which is to achieve a balance between prediction performance and model explainability. (3) DF-NAM where we replace the EBM with a neural additive model (Agarwal et al., 2021), which learns a linear combination of neural networks that each attend to a single input feature. This way, it remains explainable while being able to model non-linear effects.
4.3. Implementation details
We implemented our framework in Python 3.8. For the hate speech clustering, we use BERTopic model all-MiniLM-L6-v2 (Grootendorst, 2022) to generate embeddings and UMAP to reduce the embeddings into a seven-dimensional space. To cluster, we use HDBSCAN with a minimum cluster size of 10. We name the resulting cluster with the LLAMA-3 model Llama-3-70b-chat-hf (Meta, 2024). To model the effects, we use the EBMRegressor of InterpretML (Nori et al., 2019). We allow for ten interactions in the EBM model and run it for a maximum of 5,000 rounds. Moreover, we performed hyperparameter tuning (see Appendix B). For evaluation, we split our dataset into train/test set (80/20). To evaluate the performance of the entire framework (step  3), we use the root mean squared error (RMSE). We average the performance of the EBMRegressor as well as our baselines over five runs. To model PLV, we use BPRMF (Rendle et al., 2009). We set the learning rate of our BPRMF model to 0.001, the batch size to 64, and the embedding size to 64. We average the performance over five runs with varying train/test splits. For the IPS reweighting with $\pi^{neural}_{h}$, we employ linear discriminant analysis (LDA) topic modeling [^2] https://scikit-learn.org/stable/modules/generated/sklearn.discriminant_analysis.LinearDiscriminantAnalysis.html to get a latent representation $e_{h}$ of the topics in the posts.
5. Results
5.1. Hate speech cluster
We begin by reporting the hate speech clusters resulting from our topic modeling. We cluster the 270 hate speech root posts into four clusters, which we labeled automatically. The optimal number of clusters is determined automatically by HDBSCAN, which allows us more flexibility since we do not have to determine the number of clusters beforehand (Campello, Ricardo J. G. B. et al., 2013).
[…]
Model Explainable Flexible RMSE
DF-linear [Uncaptioned image] [Uncaptioned image] 0.0885
DF-ExNN [Uncaptioned image] [Uncaptioned image] 0.0911
DF-NAM [Uncaptioned image] [Uncaptioned image] 0.1342
0.0863
DF-EBM (ours) [Uncaptioned image] [Uncaptioned image] (--2.5%)
Table 1. Performance comparison (measured by the RMSE) for predicting users’ probability to reshare hate speech. Our debiasing framework performs best (relative improvement of 2.5% over the best baseline, whereby DF-linear is the only baseline that we are aware of in the literature).
5.2. Performance comparison
We next report the performance of our framework when predicting the effects of user attributes on the users’ probability to reshare hate speech. We compare our framework against multiple baselines based on the RMSE (root mean squared error). To the best of our knowledge, there is only one baseline in the literature: DF-linear (Cheng et al., 2021), which builds upon a linear approach and thus can not account for non-linearites. All of the baselines as well as our model are explainable since we aim to provide an understanding of the user attributes that make users reshare hate speech.
Table 1 shows the RMSE for each variant of our framework. Our debiasing framework DF-EBM outperforms the DF-linear baseline with a relative improvement of 2.5%. The improvement is statistically significant according to Welch’s $t$-test (Welch, 1947) at a $p<0.001$ level. Importantly, both frameworks have access to identical data so the performance gain must be attributed solely to the flexibility of our framework in capturing nonlinearities. Moreover, our DF-EBM also outperforms the other custom baselines that are non-linear. Here, the Welch’s $t$-tests (Welch, 1947) are again statistically significant at the $p<0.01$ level. Together, this demonstrates that our debasing framework DF-EBM is superior.
5.3. Analyzing hate speech resharing
We now leverage the explainability of our framework to understand which user attributes affect the users’ probability to reshare hate speech. For this, we analyzed the effects of users’ verified status, account age, number of posts, number of followers, and number of friends. We first look at the feature importance for each user attribute, which is computed as the average absolute predicted value of each feature for the training set (i.e., for each feature, we predict the absolute value for each sample in the training set and then average these scores). This is weighted by the relative frequency of each feature in the train set. Here, we find that the number of posts is the most influential user attribute when predicting users’ probability to reshare hate speech (see Figure 1). Moreover, the second most influential attribute is the number of followers. Thereby, we add to the literature where the popularity of a user and their activity level are important indicators as to whether they post hate speech (Hua et al., 2020; Ribeiro et al., 2018), whereas we show that the number of posts and followers also cause a larger propensity to reshare hate speech.
We next look at the effects of each user attribute on the probability to reshare hate speech. Figure 2 shows the contribution of each user attribute. The contribution is shown on the same scale as the output variable, and the graph is normalized so that the average prediction is at 0. For example, a score of $-0.1$ at an account age of 100 days would mean that a user with a 100-day-old account is 10 percentage points less likely to reshare hate speech than the average prediction. Overall, we find that users with fewer posts, fewer followers, and fewer friends are more likely to reshare hate speech. In return, newer accounts are less likely to do so. Moreover, we find that verified users are 0.016 (±$\pm 0.004$) percentage points more likely to reshare hate speech than the average prediction, which is contrary to our expectations that a verified status is typically linked to users with a large reputation who should thus engage less with hate speech. We further perform a comparison of PLV to hate speech and PLV to normal content (see Appendix E for the complete analysis).
Figure 1. Feature importance of the user attributes when predicting the users’ probability to reshare hate speech.
(a) account age
(b) #posts (in log scale)
(c) #followers (in log scale)
(d) #friends (in log scale)
Figure 2. Contributions of the user attributes to the users’ probability to reshare hate speech as estimated by our debiasing framework (both y-axis and outcome $Y$ are shown on the same scale). Each graph is normalized for better comparability such that the average prediction on the train set is at 0. The blue curve shows the predicted effect, while the orange curves show the upper and lower bounds of our model.
5.4. Resharing of types of hate speech
We now split the hate speech posts into their respective types and analyze the resharing for each type separately. Here, we look at the effects of users’ verified status, account age, number of posts, followers, and friends on the resharing behavior. Similar to above, the contribution of each user attribute is shown on the same scale as the output variable.
[…]
(a) account age
(b) #posts (in log scale)
(c) #followers (in log scale)
(d) #friends (in log scale)
Figure 3. Contributions of the user attributes to the users’ probability to reshare racist and misogynistic rants as estimated by our debiasing framework (both y-axis and outcome $Y$ are shown on the same scale). Each graph is normalized for better comparability such that the average prediction on the train set is at 0. The blue curve shows the predicted effect, while the orange curves show the upper and lower bounds of our model.
(a) account age
(b) #posts (in log scale)
(c) #followers (in log scale)
(d) #friends (in log scale)
Figure 4. Contributions of the user attributes to the users’ probability to reshare anti-Trump and anti-right wing rants as estimated by our debiasing framework (both y-axis and outcome $Y$ are shown on the same scale). Each graph is normalized for better comparability such that the average prediction on the train set is at 0. The blue curve shows the predicted effect, while the orange curves show the upper and lower bounds of our model.
(a) account age
(b) #posts (in log scale)
(c) #followers (in log scale)
(d) #friends (in log scale)
Figure 5. Contributions of the user attributes to the users’ probability to reshare toxic rants as estimated by our debiasing framework (both y-axis and outcome $Y$ are shown on the same scale). Each graph is normalized for better comparability such that the average prediction on the train set is at 0. The blue curve shows the predicted effect, while the orange curves show the upper and lower bounds of our model.
5.5. Ablation study
We performed an ablation study to analyze the differences in performance when using different inputs to the past latent vulnerability (PLV). In particular, we compare four different inputs: (1) a base model where we do not control for PLV; (2) BPRMF-V model with the virality-based PLV using $P_{\mathrm{virality}}$; (3) BPRMF-F with the follower-based PLV using $P_{\mathrm{follower}}$; and (4) BPRMF-NN with the neural network-based PLV using $P_{\mathrm{neural}}$. Table 2 shows the RMSE for the five different inputs. Overall, all models that control for PLV outperform the base model that does not control for it, although we do not see a performance gain that is statistically significant based on Welch’s $t$-test at the $p<0.05$ level. In sum, the ablation study confirms the effectiveness of IPS reweighting for modeling PLV. In the following experiments, we will use the debiasing framework with the virality-based PLV.
User
Model attributes PLV RMSE
Base [Uncaptioned image] 0.0866
BPRMF-V [Uncaptioned image] Virality-based propensity 0.0863
BPRMF-F [Uncaptioned image] Follower-based propensity 0.0864
BPRMF-NN [Uncaptioned image] NN-based propensity 0.0865
Table 2. Ablation study using different inputs for the past latent vulnerability. Results are averaged over 5 runs.
6. Discussion
Summary of findings: In our paper, we show the effects of user attributes on hate speech resharing while controlling for the past latent vulnerability of a user to hate speech. We find that users with fewer posts, fewer followers, and fewer friends tend to reshare more hate speech and newer accounts tend to reshare less. When splitting hate speech into different types, we find that racist and misogynistic rants are spread mostly by newer accounts, accounts with fewer posts, accounts with fewer friends, and accounts with a moderate number of friends. In return, political anti-Trump and anti-right-wing rants are reshared by older accounts, accounts with a moderate number of posts, and accounts with many followers and friends. Toxic rants are also shared by older accounts and accounts with a large number of posts, followers, and friends.
Interpretation: A possible explanation for our findings might be the concept of social influence. Social influence is one of the most important determinants in explaining the spread of information on social media (Zaman et al., 2014). Users with many followers tend to have more social influence, and, therefore, their posts reach a large audience (Suh et al., 2010; Pröllochs and Feuerriegel, 2023; Maarouf et al., 2024; Vosoughi et al., 2018; Zaman et al., 2014). Similarly, verified users are perceived to have more credibility online (Morris et al., 2012) and, hence, have higher social influence. Users with a large social influence (e.g. verified, older accounts, more posts, more friends, and more followers) might be more concerned about having a positive image (Talwar et al., 2019), resulting in a lower probability to reshare harmful content in general.
When looking at the different types of hate speech, we make the interesting observation that the dynamics behind the of spread racial and misogynistic hate and political hate differ. Users with little social influence tend to spread more racial and misogynistic hate while users with high social influence tend to spread political hate. This might be the case because racial and misogynistic hate is socially unacceptable and circulates primarily in small online communities where extreme views are normalized and users do not have to fear backlash for resharing. In return, the political landscape is widely discussed online and political hate might border what is still deemed acceptable to express online or is harder to detect for moderation efforts.
Relevance: Regulatory initiatives in many countries around the world (e.g., the Digital Services Act in the European Union) compel social media platforms to limit the spread of online hate speech. In this regard, our framework helps platform providers identify which user attributes make users reshare hate speech. This allows for the adoption of a user-centric design approach when developing targeted and data-driven mitigation strategies. By tailoring interventions to address the roots of hate speech resharing, we can implement more effective measures to prevent users from sharing hate speech and, thereby, counter its spread.
Implications: Our framework shows the influences of user attributes on the probability to reshare hate speech and therefore generates insights on how to take action against the resharing of hate speech. For example, online platforms might consider not showing the popularity of a user by removing follower counts from the profile information. This way, users are unaware of how many people see their posts, which might lead them to be more selective in the content they share.
Platforms could also vary their moderation measures depending on the user’s follower count, e.g., decrease the response times to reports against hate speech from popular users. However, this would require a shift in resources and could raise ethical considerations about the uneven moderation of users. Given the positive effects of a verified status on hate speech resharing, online platforms might also consider removing the verified status. This might lead to users being more responsible in the content they reshare.
Conclusion: In this work, we leverage a state-of-the-art debiasing framework to model the effects of user attributes on the probability to reshare hate speech while controlling for the past latent vulnerability to hate speech. Our framework provides flexible and explainable results while achieving state-of-the-art performance. We also cluster hate speech into its multiple types, which allows us to model the resharing dynamics individually.
Author contributions
All authors contributed to conceptualization, results interpretation, and manuscript writing. First author contributed to data analysis. All authors approved the manuscript.
Appendix A Statistics of user attributes
In our work, we estimate the effects of five user attributes on the probability to reshare: verified (= 1 if the user is verified, 0 otherwise); account age (in days); #post (number of posts by the user); #follower (number of followers of the user); and #friend (number of friends of the user). The below Table 3 shows the summary statistics of the user attributes.
User attributes Mean Median Std. dev.
verified 0.003 0 0.05
account age (in days) 3663.74 3585 833.51
#post 57588.09 24465.5 106129.73
#follower 2230.27 546 33567.04
#friend 1253.2 491 3639.85
Table 3. Summary statistics of the user attributes.
Appendix B Hyperparameter tuning
Hyperparameters are tuned via cross-validation with a 70/30 split. Table 4 shows the list of hyperparameters of DF-EBM. We fine-tuned the DF-EBM with the BPRMF-V confounder and applied the best-performing parameters to all our variants of DF-EBM model. As a result, we set the learning rate to 0.01, the number of maximum bin to 512, and the minimum sample leaf to 3.
Model Hyperparameter Search range
Learning rate 0.1, 0.01, 0.001
Maximum bins 128, 256, 512
DF-EBM Minimum sample leaf 1, 2, 3
Table 4. Hyperparameters and search ranges for DF-EBM.
Appendix C Insights into the different types of hate speech
[…]
Hate speech cluster Reshares Characteristic words
Racist and misogynistic rants 15.219 hate, niggas, nigga, dont, yall, like, ass, ugly, bitch, fucking
Anti-Trump and anti-right wing rants 6.207 idiot, trump, isis, idiots, syria, says, attack, obama, war, run
Toxic rants 512 dont, women, fucking, brutal, disgusting, life, white, just, bad, lifequotes
Social media noise 51 summer, tomorrow, selby, model, focus, views, hes, make, time, just
Table 5. Number of reshares and characteristic words of each hate speech cluster defined as the top-10 words with regard to c-TF-IDF weighting.
Hate speech cluster Example posts
[…]
Table 6. Example posts from each hate speech cluster
Appendix D Robustness checks
Modeling past latent vulnerability: We perform a robustness check to validate our choice of how we model the past latent vulnerability (PLV). Recall that we model PLV using the recommendation algorithm BPRMF. However, we have multiple ways to debias propensity scores, implying that we can use multiple inputs to the BPRMF model. Here, we check the prediction performance of the BPRMF with different propensity score inputs. We measure the prediction performance using recall@$k$ and NDCG@$k$. The former measures how many of the predicted posts were interesting to the user out of those that the model predicted to be interesting as the top-$k$ posts. The latter measures the accuracy of the prediction by comparing the ground-truth interestingness to the predicted ranking of posts for the top-$k$ ranked posts. This way, we can get an estimate of how well the recommendation algorithm was able to model PLV for each user.
Table 7 shows the prediction performance of BPRMF with different inputs. We find that the BPRMF algorithms with debiased propensity scores are better at predicting hate speech that the user is vulnerable to. This confirms our choice to employ IPS reweighting in our debiasing framework when modeling PLV.
recall@$k$
Model $k=20$ $k=40$ $k=60$ $k=80$
BPRMF 19.35 30.90 36.20 40.17
BPRMF-V 18.02 31.18 36.76 40.69
BPRMF-F 19.48 32.11* 37.92* 42.16*
BPRMF-NN 19.23 31.16 36.62 40.80
(a) Performance evaluated based on recall@$k$
NDCG@$k$
Model $k=20$ $k=40$ $k=60$ $k=80$
BPRMF 7.02 9.39 10.33 10.97
BPRMF-V 6.67 9.37 10.36 11.00
BPRMF-F 7.04 9.63 10.66 11.35*
BPRMF-NN 6.82 9.27 10.24 10.92
(b) Performance evaluated based on NDCG@$k$
Table 7. Comparison of the prediction performance (measured in recall@$k$ and NDCG@$k$) of BPRMF when modeling PLV with different propensity score inputs. An asterisk * indicates that the improvement measured with Welch’s $t$-test (Welch, 1947) is statistically significant at $p<0.05$ level.
Robustness check for smoothing operator In IPS for reweighting observational social media data, the IPS for rare events could be very high. A smoothing operator can lead to more stable propensity scores and prevent overfitting. In this work, $\mu$ is the smoothing operator. We chose $\mu=0.5$ since popularity-related measures follow a power law distribution, i.e., $\mu\leq 1$. We performed a new robustness check to study the sensitivity of the recommendation algorithm with $\mu=[0.1,0.5,1.0]$ for the virality and the follower-based propensity score. The results with respect to recall@$k$ are in Table 8. We found that the overall findings remain robust.
recall@$k$
Model $k=20$ $k=40$ $k=60$ $k=80$
BPRMF-V $\mu=0.1$ 19.41 31.47 36.40 40.50
BPRMF-V $\mu=0.5$ 18.02 31.18 36.76 40.69
BPRMF-V $\mu=1.0$ 18.56 31.08 36.72 40.53
BPRMF-F $\mu=0.1$ 18.73 30.10 35.28 39.32
BPRMF-F $\mu=0.5$ 19.48 32.11 37.92 42.16
BPRMF-F $\mu=1.0$ 18.52 31.74 37.04 41.14
Table 8. Performance comparison of BPRMF with varying degrees of $\mu$.
Appendix E Comparison of the past latent vulnerability to normal and hate speech
To see whether PLV to hate speech is different from PLV to normal content, we apply the above recommendation model to users who share normal content. We then compare the PLV embeddings of hate speech and normal content by visualizing the differences with t-SNE (van der Maaten and Hinton, 2008). In addition, we cluster the hate speech and normal content embeddings using DBSCAN (Ester et al., 1996) and compute the silhouette score. The silhouette score ranges from $-1$ to $+1$ and measures the separability of clusters.
Overall, we observed the following. In the t-SNE plot, we show that PLV embeddings of normal content are mostly concentrated in one broad cluster whereas the embeddings of hate speech are separated into multiple, clearly defined clusters (see Figure 9). We also find slightly higher silhouette scores for hate speech than normal content (0.609 and 0.539, respectively) when keeping the DBSCAN parameters constant (eps $=0.5$ and minimum samples $=10$). This suggests that PLV to hate speech is more homogenous within clusters and heterogenous across clusters, whereas PLV to normal content cannot be clustered into clearly distinct groups as well. This is in line with previous research which states that hate speech spreaders are more similar to each other (Beatty, 2020; Evkoski et al., 2022) and interact in highly connected networks (Maarouf et al., 2024; Mathew et al., 2019).
[Uncaptioned image] [Uncaptioned image]
(a) Past latent vulnerability to hate speech (Silhouette score $=0.609$) (b) Past latent vulnerability to normal content (Silhouette score $=0.539$)
Table 9. Comparison of the past latent vulnerability for (a) hate speech and (b) normal speech using a 2D $t$-SNE plot (van der Maaten and Hinton, 2008).